# Li Tana,b , Shengrong Wanga,c

### v1

@@ 4 Propagation of Chaos
$\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}),$
For $q<p$, by Lemma 4.1, we get
$\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}=$ $\mathbb{E}[(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{q}]^{p/{q}}$
$\leq$ $\mathbb{E}\bigg{(}2^{q-1}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{q}+2^{q-1}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{q}\bigg{)}^{p/q}$
$\leq$ $C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{p}$
$\leq$ $C\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j}-X_{s}^{j,N}|^{q}\bigg{)}^{p/q}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$
$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}.$
Applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|\Theta_{s}^{i}|^{p}\bigg{)}$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
According to (2.7), we obtain
$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{t-\tau}^{i}-X_{t-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Theta_{s}^{i}|^{p}.$
Noting that $10<\lambda<$ 1, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Theta_{s}^{i}|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
Since $(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$ is controlled by the Wasserstein distance estimate([13], Theorem 1), for $\overline{p}>2p$, we get
$\mathbb{E}(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right.$
Then, by Lemma 4.2, we complete the proof. ∎
5 The tamed theta scheme
In this section, we consider the tamed theta scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
@@ 5 The tamed theta scheme
for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$. Assume that for any time $T>0$, there exist positive constants $M,m\in\mathbb{N}$ such that $\Delta=\frac{\tau}{m}=\frac{T}{M}$, where $\Delta\in(0,1)$ is the step-size. Consequently, for the time grid $t_{k}=t_{0}+k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$. For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$ and $t_{0}=0$. For $k=0,1,\cdots,M$, we now define the discrete EM scheme for (4.2) as follows,
@@
where the empirical measures $\mu_{t_{k}}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{Y_{t_{k}}^{j,N}}(\cdot)$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
@@
and the continuous EM solution writes as
@@
where $\overline{\mu}_{s}^{Y,N}:=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{\overline{Y}_{t}^{j,N}}(\cdot)$, and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ when $t_{k}\in[-\tau,0]$. However, this $Y_{\Delta}^{i,N}(t)$ is not $\mathcal{F}_{t}$-adapted, it does not meet the fundamental requirement in the Itô stochastic analysis. To avoid calculus, we use the discrete split-step theta scheme (see [28]) described as follows: For $k=-m,\cdots,-1$, set $z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$. For $k=0,1,\cdots,M-1$, we rewrite (5.2) as
(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ z_{t_{k+1}}^{i,N}=D(z_{t_{k+1-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right.$
@@
$z_{t_{k+1}}^{i,N}-D(z_{t_{k+1-m}}^{i,N})=z_{t_{0}}^{i,N}-D(z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\mu_{t_{0}}^{i,N})\Delta+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta$
$+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
Similarly, we define the corresponding continuous-time split-step tamed theta solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0]$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\mu_{0}^{Y,N})\Delta$. For any $t\in[0,T]$,
@@
where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we have
@@
It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
@@
There exists a positive constant $K_{7}\geq 1$ such that
$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})),|b(x,y,\mu)|\}.$
@@
$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})).$
@@
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
@@
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2p}),$
where $K_{8}$ is a positive constant.
@@
We divide it into two case. Case 1, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)<0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}-\frac{b(\overline{x},\overline{y},\nu)}{1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Case 2, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)>0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$+\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{\Delta^{\alpha}[b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)-b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Furthermore,
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p},$
the second part is obvious. ∎
@@
Given Lemmas 5.1, 5.2 hold. We have
$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2}).$
5.1 Moment bounds

### v2

@@ 4 Propagation of Chaos
$\mathbb{W}_{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{2}(\mu_{s}^{X},\mu_{s}^{X,N}),$
then, for $p\geq 2$, Lemma 4.1 leads to
$\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})=$ $\mathbb{E}[\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})]^{p/{2}}$
$\leq$ $\mathbb{E}\bigg(2\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+2\mathbb{W}_{2}^{2}(\mu_{s}^{X},\mu_{s}^{X,N})\bigg)^{p/2}$
$\leq$ $C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+C\mathbb{E}\mathbb{W}_{2}^{p}(\mu_{s}^{X},\mu_{s}^{X,N})$
$\leq$ $C\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i}-X_{s}^{i,N}|^{2}\bigg)^{p/2}+C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})$
$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}).$
Putting $H_{1}(t)$-$H_{3}(t)$ together, we conclude that
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{p}\bigg)\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s$
$\leq$ $\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r$
$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{r}^{i}},\mu_{r}^{X}){\rm d}r.$
According to Lemma 2.7, it derives
$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Upsilon_{s}^{i}|^{p}.$
Noting that $0<\lambda<1$, then
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg)\leq$ $C\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{p}\bigg).$
That is,
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg)$
$\leq$ $\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r$
$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{r}^{i}},\mu_{r}^{X}){\rm d}r.$
Since $\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})$ is controlled by the Wasserstein distance estimate([35], Theorem 1), for $2\leq p<\overline{p}/2$, it follows
$\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right$.
Then, similar to (4.4), by Lemmas 2.8 and 4.2, the desired result can be obtained. ∎
5 The Tamed Theta EM Scheme
In this section, we consider the tamed theta EM scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
@@ 5 The Tamed Theta EM Scheme
for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$. Assume that for any time $T>0$, there exist positive constants $M,m\in\mathbb{N}$ such that $\Delta=\frac{\tau}{m}=\frac{T}{M}$, where $\Delta\in(0,1)$ is the step-size. Consequently, for the time grid $t_{k}=k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$. For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$. For $k=0,1,\cdots,M$, we now define the discrete tamed theta EM scheme for (4.2) as follows,
@@
where the empirical measures $\mu_{(\cdot)}^{Y,N}:=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{Y_{(\cdot)}^{i,N}}$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
@@
and the continuous tamed theta EM solution writes as
@@
where $\overline{\mu}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{Y}_{\Delta}^{i,N}}(\cdot)$, $\overline{\mu}^{Y_{+},N}(\cdot):=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{Y}_{\Delta+}^{i,N}}(\cdot)$ and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ for $t\in[-\tau,0]$. However, this $Y_{\Delta}^{i,N}(t)$ is not $\mathcal{F}_{t}$-adapted, it does not meet the fundamental requirement in the Itô stochastic analysis. To avoid calculus, we use the discrete split-step theta scheme (see [26]) described as follows: For $k=-m,\cdots,0$, set $Z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$. For $k=0,1,\cdots,M-1$, rewrite (5.2) as
(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+Z_{t_{k}}^{i,N}-D(Z_{t_{k-m}}^{i,N})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ Z_{t_{k+1}}^{i,N}=D(Z_{t_{k+1-m}}^{i,N})+Z_{t_{k}}^{i,N}-D(Z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right$.
@@
$Z_{t_{k+1}}^{i,N}-D(Z_{t_{k+1-m}}^{i,N})$
$=$ $Z_{t_{0}}^{i,N}-D(Z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\delta_{\xi(0)})\Delta$
$+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
Similarly, we define the corresponding continuous-time split-step tamed theta EM solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0)$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\delta_{\xi(0)})\Delta$. For any $t\in[0,T]$,
@@
where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we see
@@
It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta EM solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta EM solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
@@
There exists a positive constant $K_{6}\geq 1$ such that
$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{6}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{2}(\mu,\delta_{0})),|b(x,y,\mu)|\}$
for $x,y\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@
$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{6}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{2}(\mu,\delta_{0})).$
@@
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{7}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+\mathbb{W}_{2}^{2}(\mu,\nu)),$
@@
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+\mathbb{W}_{2}^{2p}(\mu,\delta_{0})),$
where $K_{7},K_{8}$ are positive constants.
@@
For the first inequality, we write
$b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)$
$=$ $\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|)}$
$+\frac{\Delta^{\alpha}[|b(\overline{x},\overline{y},\nu)|\cdot b(x,y,\mu)-|b(x,y,\mu)|\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|)}.$
Using Assumption 2.4 for the first term, the Lipschitz continuity of $b$ and its growth condition from Remark 2.6 along with the Young’s inequality for the second term, we obtain the desired result. Furthermore, for the second inequality, we observe that
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}$ $=\left|b(x,y,\mu)-\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}\right|^{p}$
$=\left|\frac{\Delta^{\alpha}|b(x,y,\mu)|\cdot b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}\right|^{p}$
$\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p}.$
Then, with the growth condition of $b$, the second part can be shown. ∎
@@
Let Assumption 2.4 hold. We derive directly from Lemmas 5.1, 5.2 that
$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0})).$
5.1 Moment Boundness